# The spin-one Duffin-Kemmer-Petiau equation revisited: analytical study of its structure and a careful choice of interaction

### v1

@@ III.2 $(-1)^{j}$–Parity Solution
from which the general expression of the eigenenergies is deduced, which can be rewritten as
@@
Let us now consider the unnatural parity solutions.
@@ III.3.1 $H_{0}(r)$ Solutions
$\alpha=E_{n}^{2}-M^{2}-z_{r}^{2},\quad\beta=-2x_{r}z_{r},\quad\omega=-\left(j(j+1)+x_{r}(x_{r}+1)+2y_{r}z_{r}\right)\quad\delta=-2(x_{r}+1)y_{r},\quad\sigma=-y_{r}^{2},$ (24)
equations from which the possible values of the energy are deduced:
@@ III.3.2 $\phi(r)$ Solutions
which give the energy relation as
@@ V Conclusion
Aknowledgment
The research of L.M.N. and S.Z. was supported by the Q-CAYLE project, funded by the European Union-Next Generation UE/MICIU/Plan de Recuperacion, Transformacion y Resiliencia/Junta de Castilla y Leon (PRTRC17.11), and also by RED2022-134301-T, PID2020-113406GB-I00 and PID2023-148409NB-I00, financed by MICIU/AEI/10.13039/501100011033. The work of M.B. was supported by the Czech Science Foundation within the project 22-18739S.

### v2

@@ III.2 $(-1)^{j}$–Parity Solution
from which, together with (13), the general expression of the eigenenergies is deduced, which can be rewritten as
@@
Moreover, note that the potential parameters in (21) and (22) must fulfill the constraint given by (19). Let us now consider the unnatural parity solutions.
@@ III.3.1 $H_{0}(r)$ Solutions
$\alpha=E_{n}^{2}-M^{2}-z_{r}^{2},\quad\beta=-2x_{r}z_{r},\quad$ (24)
$\omega=-\left(j(j+1)+x_{r}(x_{r}+1)+2y_{r}z_{r}\right)\quad$
$\delta=-2(x_{r}+1)y_{r},\quad\sigma=-y_{r}^{2},$
from which, together with (13), the possible values of the energy are deduced:
@@
Recall that the potential parameters in both relations (26) and (27) must satisfy the constraint given in (19).
@@ III.3.2 $\phi(r)$ Solutions
which, together with (13), give the energy relation as
@@
Note that, as in the previous cases, the potential parameters in Eqs. (29) and (30) must satisfy the constraint given in (19).
@@ V Conclusion
Acknowledgments
The work of M.B. was supported by the Czech Science Foundation within the project 22-18739S. The research of L.M.N. and S.Z. was supported by the Q-CAYLE project, funded by the European Union-Next Generation UE/MICIU/Plan de Recuperacion, Transformacion y Resiliencia/Junta de Castilla y Leon (PRTRC17.11), and also by RED2022-134301-T and PID2023-148409NB-I00, financed by MI-CIU/AEI/10.13039/501100011033. Financial support of the Department of Education of the Junta de Castilla y León and FEDER Funds is also gratefully acknowledged (Reference: CLU-2023-1-05).